OpenStudy (anonymous):

. Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. Provide the solution to the system and explain what the solution represents on the graph. 3x + 6y = 18 2x + 5y = 20

OpenStudy (anonymous):

Graph the two lines: 3x + 6y = 18 2x + 5y = 20 Each line represents all coordinates x and y that satisfy the corresponding equation. i.e. the first line consists of points that satisfy 3x + 6y = 18 Where the two lines intersect, is the value for x and the value for y that satisfy both equations. If the lines don't intersect, then no x and y satisfy both equations. If the lines are on top of each other, then all values for x and y satisfy both equations